OpenStudy (anonymous):

Consider f(x) = e^(7x − 3) Find the inverse of the function. f^ (−1)(x) = Would you be able to show me the steps to solve the problem? Thanks.

zepdrix (zepdrix):

\[\Large f(x)=e^{7x-3}\]Writing f(x) as y,\[\Large y=e^{7x-3}\]Swapping x and y,\[\Large x=e^{7y-3}\] This new \(\Large y\) in the problem represents \(\Large f^{-1}(x)\), so we want to try and solve for it.

zepdrix (zepdrix):

The next step is a little tricky. We want to take the natural log of both sides.\[\Large \ln(x)=\ln\left(e^{7y-3}\right)\]From there we can apply a rule of logarithms:\[\Large \color{royalblue}{\log(a^b) \quad=\quad b\cdot\log(a)}\]

zepdrix (zepdrix):

Understand what that will do on the right side?

OpenStudy (anonymous):

Yes, thank you.

OpenStudy (anonymous):

lnx = 7y-(3lne) then 1/7(lnx) = 7/7-((3lne))/7 Am I right so far?

zepdrix (zepdrix):

Hmm your step where you pulled the exponent out of the log looks a little weird. We want to pull it down in it's own set of brackets, like this:\[\Large \ln(e^{7x-3}) \quad=\quad (7x-3)\ln(e)\]

zepdrix (zepdrix):

oh that's a y, not an x. my bad :)

OpenStudy (anonymous):

(7y-3)lne = lnx then what? lne =lnx^(7y-3)

OpenStudy (anonymous):

correction lnx =lny^(7y-3)

zepdrix (zepdrix):

From here:\[\Large \ln(e^{7x-3}) \quad=\quad (7x-3)\color{royalblue}{\ln(e)}\] We would make note of this fact ~ If we take the log of a number which matches the base, then the result is simply 1. Examples: \[\Large \log_2(2)=1\]\[\Large \log_e(e)=1\]\[\Large \log_{10}(10)=1\]

zepdrix (zepdrix):

Ahhh I meant to type,\[\Large \ln x \quad=\quad (7x-3)\color{royalblue}{\ln(e)}\]on the first line. anyway,

zepdrix (zepdrix):

So using that information, and remembering what the natural log actually is,\[\Large \ln(e) \quad=\quad \log_e(e) \quad = \quad ?\]

zepdrix (zepdrix):

Wow i keep screwing up the variables.. sorry about that :(\[\Large \ln x \quad=\quad (7y-3)\color{royalblue}{\ln(e)}\]